# CogDDN: A Cognitive Demand-Driven Navigation with Decision Optimization and Dual-Process Thinking

### v1

@@ 3.4. Heuristic Process
| Seen Ins. | Unseen Ins. | Seen Ins. | Unseen Ins. | |||||||||
@@
| CogDDN(ours) | 38.3 | 17.2 | 29.8 | 37.5 | 18.0 | 28.6 | 33.3 | 16.4 | 27.1 | 34.5 | 17.1 | 27.5 |
3.5. Analytic Process
The Analytic Process is employed to reflect on navigation obstacles, as illustrated in Figure 6. Through comprehensive pre-training on diverse datasets, VLMs naturally accumulate broad world knowledge, enabling them to address complex challenges with sophisticated reasoning and insight. This capability aligns with the objectives of the Analytic Process, which relies on detailed analysis and contextual understanding to draw sound inferences in indoor navigation.
Specifically, in a closed-loop indoor navigation scenario where the Demand Matching and Heuristic Process modules operate, any obstruction encountered during navigation will trigger the reflection mechanism. The Analytic Process utilizes the vast world knowledge embedded within the VLMs to analyze the situation based on previously gathered information. During this process, the object location, scene description, reasoning, and decision from the preceding frame before the obstruction are forwarded to the Analytic Process. The system then carefully analyzes the cause of the event, identifies the error, and generates corrected reasoning and decisions. The insights gained from this reflection process are subsequently integrated into the knowledge base, enabling the system to learn from failures continuously. This iterative learning process progressively enhances the system’s decision-making capabilities, resulting in more informed and accurate navigation strategies for future tasks.
@@ 4.1. Data preparation
The experiments use the AI2Thor (Kolve et al., 2017) and the ProcThor dataset (Deitke et al., 2022). Following DDN (Wang et al., 2023), we evaluate 600 scenes (200 each from ProcThor’s train, validation, and test sets). These environments contain 109 distinct object categories, each capable of fulfilling the demand instructions. For instructions, we use the 1692 training, 241 validation, and 485 test instructions provided by DDN (Wang et al., 2023).
Data for 3D Robot Perception. We used data from the AI2Thor simulator to train the 3D Robot Perception Module. Specifically, we collected 55K frames, ensuring that 70% of the data was allocated for training, 10% for validation, and 20% for testing. For clarity, we provide details and methods about data in Appendix A.1.
Data for Demand Matching. By leveraging the instruction data, we curated an instruction-attribute dataset for the Demand Matching Module. The dataset was constructed by inputting instructions and associated objects into the LLM, which generated common attributes shared by objects fulfilling the instructions. We collected 7.2K data for training, 1.4K for validation, and 2.1K for testing. Additional details are provided in Appendix A.2.
Data for Heuristic Process. To support the supervised fine-tuning of our Heuristic Process using the knowledge base, we curated datasets by collecting data from the AI2Thor simulator. The construction of the knowledge base is discussed in Section 3.3. We generated approximately 72K trajectories, with 80% used for training and 20% reserved for validation. Additional details are provided in Appendix A.3.
@@ 4.2. Implementation Details
We employ UniMODE (Li et al., 2024) for 3D robot perception and GPT-4 for the Analytic Process to enable rational and logical thinking. Qwen2-VL-7B (Wang et al., 2024a) is employed for both the Exploit module in the Heuristic Process and Demand Matching, with multiple LoRA configurations for optimal performance. All experiments are trained on 4 RTX6000 GPUs. Please refer to Appendix D for further implementation details.
@@ 4.3. Evaluation in Closed-Loop Navigation
We conduct closed-loop experiments in AI2Thor, a widely-used open-source simulator, to evaluate the performance of CogDDN on 400 scenes using three metrics: Navigation Success Rate (NSR), Navigation Success Rate weighted by the path length (SPL) (Anderson et al., 2018), and Selection Success Rate (SSR). Table 2 compares our method with competitive methods offered by DDN (Wang et al., 2023) on the ProcThor benchmark.
@@ 4.4. Ablation Study
We conduct ablation studies on Exploit, CoT, and Reflection in a closed-loop navigation system. The experiments were carried out on both unseen scenes and unseen instructions, demonstrating the generalization and continuous learning capabilities of our CogDDN.
@@
Ablation study of Reflection. Reflection plays a crucial role in enhancing the continuous learning capabilities of our proposed CogDDN. It involves analyzing mistakes and integrating correct experiences into the knowledge base, thereby promoting self-improvement through proactive summarization and the accumulation of experiences in previously unseen environments. In our closed-loop experiment, we conducted four rounds of 500 epochs each, where the experiences generated by the reflection mechanism were added to the knowledge base and used to fine-tune the VLM. As shown in Figure 7, after four rounds of iterations, both SSR and NSR showed marginal improvements, as the reflection mechanism allows the agent to navigate around obstacles once they are encountered, resulting in only modest improvements in these metrics. However, SPL exhibits a significant increase, as the agent, learning from experiences, anticipates obstacles and takes proactive actions. This highlights the continuous learning capability of our CogDDN, where performance improves with experience.
@@ 5. Conclusion
In this paper, we present CogDDN, a dual-process closed-loop navigation system that mimics the human attentional mechanism by selectively focusing on key objects necessary for fulfilling user demands. We have developed a dual-process decision-making module that simulates human cognitive processes. Chain of Thought (CoT) reasoning is incorporated to enhance the decision-making process. Moreover, we have integrated a cumulative knowledge base to enable the continuous self-improvement of the algorithm. Our evaluations demonstrate that the proposed method achieves remarkable closed-loop performance while requiring limited training data.
@@ A.1. Annotations for 3D Robot Perception
We utilized data collected from the AI2THOR simulator to train the 3D Robot Perception Module. The process of obtaining 3D bounding box information involved generating point clouds for each object using segmentation and depth maps provided by the AI2THOR simulator. These point clouds were used to extract the minimal bounding box that fully encloses each object, allowing us to derive accurate 3D bounding box information.
@@ A.2. Annotations for Demand Matching
In this study, we curated a specialized instruction-attribute dataset to train the Demand Matching Module. The dataset construction process involved leveraging the provided instruction data, which was paired with the objects that could potentially fulfill the instructions.
The process begins with each instruction and its associated objects being input into a Large Language Model (LLM), which generates the common attributes shared by objects that could satisfy the instruction. These attributes are crucial for matching objects with the instruction. A QA-format dataset is then created, where the first part consists of the common attributes that define the correct object, assisting the Demand Matching Module in identifying the appropriate object. The second part includes a list of objects, selected from a predefined set, that align with the generated attributes and are deemed to fulfill the instruction. A case of Demand Matching dataset is shown in Figure 8.
@@ A.3. Annotations for Heuristic Process
To support the supervised fine-tuning of our Heuristic Process, we created datasets using the AI2THOR simulator. For each scenario, we randomly select an instruction and target object, then compute an optimal path using the A* algorithm, which is discretized into a sequence of actions guiding the agent to the target. After each action, the simulator produces updated images, used to identify the target object, enabling precise localization. When the agent detects the target, VLM generates a structured scene description and reasoning based on the target’s spatial location and navigation decisions.
As shown in Figure 9, the data is formatted as a Visual Question Answering (VQA) structure, with a question about the information of target object and an answer containing the scene description, reasoning steps, and final decision. The structured data is stored in the knowledge base, allowing the system to build a robust understanding of the scene and make informed decisions in future scenarios.
@@ C.3. Cases of reflection mechanism
Appendix D Implementation Details
We train the UniMODE model with a batch size of 64 for 30,000 iterations on 4 RTX6000 GPUs, taking approximately 10 hours. The input image resolution is 448 × 448 pixels, resized to 640 × 512 pixels. We use the AdamW optimizer with a learning rate of , weight decay of , and a CosineAnnealing scheduler with a warm-up factor of 0.3333 for the first 2000 iterations. Gradient clipping is applied at 35.0 for stability.
For supervised fine-tuning (SFT), we use the Qwen2-VL-7B model with LoRA, training with a batch size of 1 for five epochs on 4 RTX6000 GPUs in 1 hour. Key hyperparameters include a learning rate of , gradient accumulation of 8, and a cosine scheduler. LoRA is applied with a rank of 64 and alpha of 128, while freezing the ViT layers.
For the Heuristic Process, we conduct SFT on Qwen2-VL-7B for five epochs using knowledge base samples, which takes about 14 hours, with consistent hyperparameters from the Demand Matching module.

### v2

@@ 3.5. Analytic Process
| Seen Ins | Unseen Ins | Seen Ins | Unseen Ins | |||||||||
@@
[…]
@@ 4.1. Data preparation
Our experiments are conducted in the AI2Thor (Kolve et al., 2017) simulator using the ProcThor dataset (Deitke et al., 2022). Following the setup of DDN (Wang et al., 2023), we evaluate our model on 600 scenes (200 from each of ProcThor’s train, validation, and test splits), containing 109 object categories. We use the instruction sets provided by DDN (Wang et al., 2023), comprising 1692 training, 241 validation, and 485 test instructions.
Data for 3D Robot Perception. For the 3D Robot Perception Module, we collected 55K frames from AI2THOR, split into 70% training, 10% validation, and 20% testing sets. Each frame includes annotations for object category, 2D bounding box, and 3D box attributes. The 3D box information was obtained by extracting the minimal bounding box from the object’s point cloud, which was generated using the simulator’s depth and segmentation maps. For clarity, we provide details and methods about data in Appendix A.1.
Data for Demand Matching. For the Demand Matching Module, we created a dataset of 10.7K instruction-attribute pairs. The construction process involved using LLMs to generate common attributes for objects associated with a given instruction. This was structured into a QA format, where the question (Q) was an instruction and a set of candidate objects, and the answer (A) was the shared attributes and a list of correct objects. The dataset was split into 7.2K training, 1.4K validation, and 2.1K test samples. Additional details are shown in Appendix A.2.
Data for Heuristic Process. For the Heuristic Process, we generated a dataset of 72K expert trajectories in AI2THOR. An optimal path was computed for each instruction using the A* algorithm (Hart et al., 1968). Once the target object was detected in the agent’s field of view, we used a VLM to generate a rationale for the path taken. This data was formatted into a VQA structure, where the question (Q) was the target’s location and the answer (A) was the scene description, reasoning, and decision sequence. This dataset, used to populate our Knowledge Base, was split into 80% for training and 20% for validation. Additional details are provided in Appendix A.3.
@@ 4.2. Implementation Details
For the 3D Robot Perception module, we utilize UniMODE (Li et al., 2024). The model is trained for 30,000 iterations with a batch size of 64 on four RTX6000 GPUs, a process that takes approximately 10 hours. Input images are resized from 448 × 448 to 640 × 512 pixels. We use the AdamW optimizer with a learning rate of , and a weight decay of . A CosineAnnealing scheduler is used, with a warm-up factor of 0.3333 for the first 2000 iterations. To ensure stability, gradient clipping is applied at a threshold of 35.0.
For the Demand Matching, we conduct SFT on Qwen2-VL-7B (Wang et al., 2024a) with LoRA, training with a batch size of 1 for 5 epochs on 4 RTX6000 GPUs in 1 hour. Key hyperparameters include a learning rate of , gradient accumulation of 8, and a cosine scheduler. LoRA is applied with a rank of 64 and alpha of 128, while freezing the ViT layers.
For the Heuristic Process, we conduct SFT on Qwen2-VL-7B for 5 epochs using knowledge base samples, which takes about 14 hours, with consistent hyperparameters from the Demand Matching.
@@ 4.3. Evaluation in Closed-Loop Navigation
We conduct closed-loop experiments in AI2Thor, a widely-used open-source simulator, to evaluate the performance of CogDDN on 400 scenes using three metrics: Navigation Success Rate (NSR), Navigation Success Rate weighted by the path length (SPL) (Anderson et al., 2018), and Selection Success Rate (SSR). Table 2 compares our method with competitive methods offered by DDN (Wang et al., 2023) on the ProcThor.
@@ 4.4. Ablation Study
We conduct ablation studies on Exploit CoT, and Reflection in a closed-loop navigation system. The experiments were carried out on both unseen scenes and unseen instructions, demonstrating the generalization and continuous learning capabilities of our CogDDN.
@@
Ablation study of Reflection. Reflection is crucial for enhancing the continuous learning capabilities of our proposed CogDDN. This mechanism promotes self-improvement by analyzing navigational failures and integrating corrected experiences into the knowledge base, allowing the agent to proactively accumulate knowledge in unseen environments. We conducted a closed-loop experiment over four iterative rounds of 500 epochs each to evaluate this capability. In each round, experiences generated by the reflection mechanism were used to augment the knowledge base and fine-tune the VLM. As shown in Figure 7, after four rounds of iterative learning, both SSR and NSR showed only marginal gains, which is an expected outcome as the reflection mechanism primarily enables the agent to recover from obstacles after they are encountered, leading to only modest improvements in these goal-completion metrics. In contrast, the SPL exhibited a significant increase. This substantial improvement in SPL demonstrates that as the agent accumulates experience, it learns to anticipate potential obstacles and take proactive detours, resulting in more efficient paths. This outcome underscores the effective continuous learning capability of CogDDN, as its performance demonstrably improves with experience.
@@ 5. Conclusion
In this paper, we introduced CogDDN, a novel dual-process closed-loop navigation system designed to emulate human attentional mechanisms. Central to our approach is a dual-process decision-making module that simulates human cognition, where a deliberative process is enhanced by Chain of Thought (CoT) reasoning to navigate complex scenarios. Furthermore, we integrated a cumulative knowledge base that enables the agent to continuously learn and self-improve from its experiences. Our comprehensive evaluations demonstrate that CogDDN achieves state-of-the-art performance in closed-loop navigation tasks while requiring minimal training data, underscoring its efficiency and practical potential.
@@ A.1. Annotations for 3D Robot Perception
The 3D Robot Perception Module was trained on data from the AI2THOR simulator. To derive the 3D bounding box ground truth, we first generated a point cloud for each object using the simulator’s segmentation and depth maps. Subsequently, the minimal bounding box enclosing this point cloud was extracted to serve as the final, accurate 3D label.
@@ A.2. Annotations for Demand Matching
To train the Demand Matching Module, we constructed a specialized instruction-attribute dataset. This dataset was built from instruction-object pairs, where each instruction is associated with a set of candidate objects capable of fulfilling its requirements.
The construction process leverages a Large Language Model (LLM) to extract semantic features. For each instruction-object pair, the LLM is tasked with inferring the common attributes shared by the candidate objects that satisfy the instruction. These inferred attributes are then formulated into a Question-Answering (QA) format. The ”question” component consists of the common attributes defining the target object, while the corresponding ”answer” component contains a list of objects from the candidate set that match these attributes. A sample entry from the Demand Matching dataset is illustrated in Figure 8.
@@ A.3. Annotations for Heuristic Process
To generate a dataset for the supervised fine-tuning of our Heuristic Process, we utilized the AI2THOR simulator. Each scenario begins with a randomly selected instruction and a corresponding target object. We then compute an optimal trajectory to the target using the A* algorithm, which yields a discrete sequence of actions. As the agent executes this trajectory, it receives updated visual observations from the simulator at each step, enabling progressive localization of the target.
Upon successful detection of the object, a Vision-Language Model (VLM) is prompted to generate a structured output. This output includes a detailed scene description, a chain of reasoning that rationalizes the navigational choices, and the final decision. This information is then formatted into a Visual Question Answering (VQA) structure, as illustrated in Figure 9. In this format, a query regarding the target object’s status serves as the ”question,” while the VLM’s generated text constitutes the ”answer.” These structured entries populate a knowledge base, providing the robust data required for our system to learn complex scene understanding and decision-making.